OpenStudy (anonymous):

What is the solution to the system of equations? 4x+y=2 x-y=3 (_,_) Please Help

OpenStudy (gorv):

x-y=3 from here find x in term of y

OpenStudy (gorv):

add y on both side

OpenStudy (gorv):

and tell me what u got

OpenStudy (anonymous):

okay

OpenStudy (gorv):

after finding the value of x in other equation replace x by that value

OpenStudy (anonymous):

4x=2y

OpenStudy (gorv):

in x-y =2 add y on both side in this

OpenStudy (anonymous):

x=2y????

OpenStudy (gorv):

x-y=3

OpenStudy (gorv):

now put this value of x in other equation

OpenStudy (anonymous):

so it would be 3=2y

OpenStudy (anonymous):

????

OpenStudy (gorv):

how u got that??

OpenStudy (gorv):

x=3+y we got replace x 4(3+y)+y=2

OpenStudy (anonymous):

I added the 3 and got that

OpenStudy (gorv):

now distribute the 4 inside

OpenStudy (anonymous):

12+4y+y=2

OpenStudy (gorv):

yeah gud gud ....now add term with y

OpenStudy (gorv):

4y+Y=??

OpenStudy (anonymous):

so add y

OpenStudy (gorv):

yeah just forget y as both term have y 4+1=??

OpenStudy (anonymous):

5

OpenStudy (gorv):

now attach y to it so 5y

OpenStudy (gorv):

12+4y+y=2 12+5y=2

OpenStudy (gorv):

now subtract 12 from both side

OpenStudy (anonymous):

5y=-10

OpenStudy (anonymous):

is that correct

OpenStudy (gorv):

yeah u r gud

OpenStudy (gorv):

now divide both side by 5

OpenStudy (anonymous):

5=-2

OpenStudy (gorv):

\[\frac{ 5y }{ 5 }=\frac{ -10 }{ 5 }\]

OpenStudy (gorv):

on left side only 5 will get cancel by 5

OpenStudy (anonymous):

oh so it is-2

OpenStudy (gorv):

yeah y=-2

OpenStudy (gorv):

now put this value in x-y=3

OpenStudy (anonymous):

OK SO x--2=3

OpenStudy (anonymous):

is x -5

OpenStudy (gorv):

when 2 minus sigh comes together --=+ x+2=3

OpenStudy (gorv):

now subtacrt 2 from both side

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

Can you help me with another one?

OpenStudy (gorv):

post another Q seprately:)

OpenStudy (anonymous):

OK THANK YOU